OpenStudy (anonymous):

pythagorean theroem help x^2+(2x+4)^2=(2x+6)^2

OpenStudy (anonymous):

A^2+B^2=C^2

OpenStudy (anonymous):

can u help me solve it

OpenStudy (anonymous):

sure one sec...

OpenStudy (anonymous):

a=x b=2x+4 c=2x+6

OpenStudy (anonymous):

x^2+4x^2+16x+16x=4x^2+24x+36